OpenStudy (anonymous):

An airline allows each passenger 36 kilograms of luggage. Carl has to lower the weight of his suitcase by 25% to stay within the limit. How much did the suitcase originally weigh?

OpenStudy (whpalmer4):

Let M be the maximum weight allowed, and W be the original weight. We know that Carl had to remove 25% of W to make his suitcase's weight be M, and we know that M is 36 kg. (W-25% of W) = M = 36 kg But 25% is just 25/100, so we can write \[W - \frac{25}{100}W = 36 kg\]multiply both sides by 100 to get rid of the fraction\[100W - 25W = 3600 kg\]\[75W=3600kg\]Divide both sides by 75 to find W

OpenStudy (anonymous):

48?